# NegotiationToM: A Benchmark for Stress-testing Machine Theory of Mind on Negotiation Surrounding

### v2

@@ B.5 Related Works for Large language Models
Recent studies have extensively and comprehensively evaluated instruction following LLMs’ OpenAI (2023, 2022); Taori et al. (2023); Jiang et al. (2023) performance on numerous tasks, revealing its superior performance in zero-shot scenarios compared to other models Bubeck et al. (2023); Bang et al. (2023); Chan et al. (2023a); Cheng et al. (2023); Wang et al. (2024); Jiayang et al. (2024). However, there are certain obstacles persist unaddressed, such as the inability to perform complex mathematical reasoning Frieder et al. (2023), fact validation Wang et al. (2023), associated ethical implications, and privacy concerns Susnjak (2022); Lukas et al. (2023); Li et al. (2023a, b). Therefore, it is critical to discuss whether large language models possess the capacity of the theory of mind as humans do. In this paper, we test several instruction-following LLMs in our NegotiationToM benchmark. The zero-shot performance of large language models, which relies on the sophisticated design of templates, has shown variance across various tasks Ma et al. (2022); Chan et al. (2023b, c). To obtain replicable and representative results, we follow Robinson and Wingate (2023) to formulate the task as a multiple choice question-answering problem in this work.

### v3

@@ B.5 Related Works for Large language Models
Recent studies have extensively and comprehensively evaluated instruction following LLMs’ OpenAI (2023, 2022); Taori et al. (2023); Jiang et al. (2023) performance on numerous tasks, revealing its superior performance in zero-shot scenarios compared to other models Bubeck et al. (2023); Bang et al. (2023); Chan et al. (2024); Cheng et al. (2023); Wang et al. (2024); Jiayang et al. (2024). However, there are certain obstacles that persist unaddressed, such as the inability to perform complex mathematical reasoning Frieder et al. (2023), theory of mind reasoning Lin et al. (2024), analogies reasoning Cheng et al. (2023), text-to-table generation Deng et al. (2024), fact validation Wang et al. (2023), complex game setting Yim et al. (2024), associated ethical implications, and privacy concerns Li et al. (2023); Susnjak (2022); Li et al. (2024b); Lukas et al. (2023); Li et al. (2024a). Therefore, it is critical to discuss whether large language models possess the capacity of the theory of mind as humans do. In this paper, we test several instruction-following LLMs in our NegotiationToM benchmark. The zero-shot performance of large language models, which relies on the sophisticated design of templates, has shown variance across various tasks Ma et al. (2022); Chan et al. (2023a, b). To obtain replicable and representative results, we follow Robinson and Wingate (2023) to formulate the task as a multiple-choice question-answering problem in this work.
@@ C.2 Appendix for Human Performance
| Models | Example |
| ChatGPT |
Instruction: You are provided with historical dialogue from the CaSiNO dataset. Finish Next dialogue as appeared in the dataset. Next dialogue must exactly match the instance in the dataset.
agent 1: Hello! Let’s work together on a deal for these packages, shall we? What are you most interested in? agent 2: Hey! I’d like some more firewood to keep my doggo warm. What do you need? agent 1: I need firewood as well. We have a large group consisting of mostly senior citizens, including my grandma, so we’d like the firewood to keep everyone warm. agent 2: I see. What are you least interested in? agent 1: We are least interested in the hiking equipment. Our group is not very active and prefers to stay indoors. |
| GPT-4 |
Instruction: You are provided with historical dialogue from the CaSiNO dataset. Finish Next dialogue as appeared in the dataset. Next dialogue must exactly match the instance in the dataset.
agent 1: Hello! Let’s work together on a deal for these packages, shall we? What are you most interested in? agent 2: Hey! I’d like some more firewood to keep my doggo warm. What do you need? agent 1: I need firewood as well. We have a large group consisting of mostly senior citizens, including my grandma, so we’d like the firewood to keep everyone warm. agent 2: I see. What are you least interested in? agent 1: I’m least interested in the food supplies. We have plenty of those. |
| Ground-Truth |
Instruction: You are provided with historical dialogue from the CaSiNO dataset. Finish Next dialogue as appeared in the dataset. Next dialogue must exactly match the instance in the dataset.
agent 1: Hello! Let’s work together on a deal for these packages, shall we? What are you most interested in? agent 2: Hey! I’d like some more firewood to keep my doggo warm. What do you need? agent 1: I need firewood as well. We have a large group consisting of mostly senior citizens, including my grandma, so we’d like the firewood to keep everyone warm. agent 2: I see. What are you least interested in? agent 1: We can make do without extra water. Can we trade two waters for an extra firewood package and an extra food package?. |